OpenStudy (anonymous):

The Sears Tower is 225 meters tall plus one-half of its height. How tall is it? I know this is simple, but I just can't remember how to write out the initial question, I was always bad at story problems. I started with. 225 1 --- + - 1 2

OpenStudy (zarkon):

\[225+\frac{1}{2}225\]

hero (hero):

What Zarkon said...

OpenStudy (anonymous):

So if my math is correct it would be 337.5 meters? Question: How did you do the 1/2 as a fraction on here? I can't find the symbols, lol.

OpenStudy (zarkon):

\frac{1}{2}

hero (hero):

Of course, you would have to open equation editor first in order for that to work.

OpenStudy (zarkon):

just put the above between this \[

OpenStudy (zarkon):

and this \]

hero (hero):

\[\frac{1}{2} \]

OpenStudy (anonymous):

\[225 + \frac{1}{2}225 : 225 + \frac{1}{2}(\frac{225}{1}) : 225 + \frac{225}{2} : \frac{225}{1} + \frac{225}{2} : \frac{225(2)}{1(2)} + \frac{225}{2} : \frac{450}{2} + \frac{225}{2} : \frac{675}{2} = 335.5\]

OpenStudy (anonymous):

So.. is that right? The end got cut off, it should be = 337.5. The : in the equation line denotes going down a line to start the next step of the problem.

OpenStudy (zarkon):

\[225+\frac{1}{2}225=225(1+.5)=225\cdot 1.5=337.5\]

OpenStudy (anonymous):

Alright, thanks :)